# Dynamic Multi-Agent System with Stable Maneuvering for Robust GAIA Problem Solving by AWorld

### v2

@@ Abstract
The rapid advancement of large language models (LLMs) has empowered intelligent agents to leverage diverse external tools for solving complex real-world problems. However, as agents increasingly depend on multiple tools, they encounter new challenges: extended contexts from disparate sources and noisy or irrelevant tool outputs can undermine system reliability and accuracy. These challenges underscore the necessity for enhanced stability in agent-based systems. To address this, we introduce dynamic supervision and maneuvering mechanisms, constructing a robust and dynamic Multi-Agent System (MAS) architecture within the AWorld framework. In our approach, the Execution Agent invokes the Guard Agent at critical steps to verify and correct the reasoning process, effectively reducing errors arising from noise and bolstering problem-solving robustness. Extensive experiments on the GAIA test dataset reveal that our dynamic maneuvering mechanism significantly improves both the effectiveness and stability of solutions, outperforming single-agent system (SAS) and standard tool-augmented systems. As a result, our dynamic MAS system achieved first place among open-source projects on the prestigious GAIA leaderboard. These findings highlight the practical value of collaborative agent roles in developing more reliable and trustworthy intelligent systems.
Keywords Dynamic Multi-Agent System  $\cdot$ Runtime  $\cdot$ Stability  $\cdot$ Maneuvering
@@ 1 Introduction
The extraordinary progress of Large Language Models (LLMs)—in terms of both capabilities and scale (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; The Google DeepMind Team, 2024; Anthropic, 2025)—has sparked widespread curiosity about the upper bounds of artificial intelligence. As practitioners continue to push the frontiers, it has become clear that augmenting foundational models with external tools not only expands their problem-solving abilities beyond intrinsic knowledge, but also enables the tackle of complex real-world challenges  (Kapoor et al., 2024; Huang and Yang, 2025; Krishnan, 2025; Shao et al., 2025). A vivid illustration is provided by the recent surge in interest surrounding the IMO2025 mathematics competition, where state-of-the-art LLMs, even those with impressive baseline performance, struggled to solve many problems. In contrast, agent-based systems—built upon these same foundational models—successfully solved five out of six competition tasks (Huang and Yang, 2025), prompting new questions about the true limits of AI: perhaps it is not just the power or size of individual models, but also our ingenuity in organizing them for effective collaboration that also matters.
@@
This insight has fueled the rapid growth of agent frameworks, where multiple models or agents interact and work together, often supported by standardized tool interfaces like MCP (Hou et al., 2025; Liu et al., 2025). The marriage of LLMs and tools has inspired developers, but amid the excitement, a central challenge has come to the fore: system stability. Empirical results show that an agent’s robustness often hinges on the foundational model’s reliability (Coletta et al., 2024; Li et al., 2025; Shojaee et al., 2025), the nature of integrated tools, and design strategies for agent orchestration. Thus, to advance reliable intelligent systems, research must go beyond scaling up model capacities and also focus on strategies for building agents that are consistent, resilient, and adaptive.
The multi-agent paradigm, especially in forms inspired by the “solver–reviewer” structure seen in recent IMO2025 studies (Huang and Yang, 2025), offers a simple yet effective blueprint: two agents, one solving and the other reviewing, boost overall problem-solving accuracy compared to single-agent setups. Yet, such enforced dialog-based architectures introduce longer contexts and extended reasoning chains, with practical limitations such as rigid turn limits for ending conversations. This points to the desirability of more flexible, dynamic multi-agent systems.
Drawing inspiration from disciplines like vessel maneuvering—where optimal navigation depends on dynamically adjusting control parameters based on hydrodynamics conditions rather than static settings (Xie et al., 2020)—we propose a parallel in agent orchestration. Intelligent agents should not rely on fixed, predetermined supervision mechanisms. Instead, they require dynamic maneuvering: agents ought to adaptively decide when and how to intervene or correct reasoning processes, based on the evolving task context, tool outputs, and model-internal interpretations.
@@
Motivated by these challenges and opportunities, this work introduces a dynamic maneuvering mechanism within the AWorld framework, our open-source agent development platform. By enabling a principal Execution Agent to collaborate with a Guard Agent that verifies and refines the reasoning process at critical moments, we construct a flexible multi-agent system. Rigorous testing on the GAIA benchmark—renowned for assessing agent capabilities—demonstrates (Mialon et al., 2023) that our dynamically orchestrated MAS not only surpasses single-agent systems (SAS) built on the same foundation and tools, but also achieves higher stability and accuracy. As a result, our MAS system achieved the first place among open-source projects on the prestigious GAIA test leaderboard, as shown in Figure 2. This research highlights the promise of adaptive multi-agent approaches for building powerful, trustworthy intelligent systems equipped to face diverse real-world tasks.
@@ 2 Method
Our approach to enhancing agent stability is inspired by well-established principles in control theory, particularly from the field of marine vessel navigation (Xie et al., 2020). The maneuvering motions of a ship, as described by the control equations presented in Table (1), model a system subject to complex and unpredictable external forces such as wave drift. To ensure the vessel converges to a desired trajectory (for example the zig-zag test as shown in Figure 3), continuous and adaptive adjustments are essential. The rudder angle, for instance, is dynamically controlled to counteract disturbances and guide the ship’s path.
We posit that an agent’s problem-solving process is analogous to a ship’s voyage. The agent’s logical reasoning forms a "trajectory" toward a solution, which is constantly perturbed by "noise" from long contexts and irrelevant tool outputs—akin to the winds and currents affecting a vessel. Drawing from this parallel, we introduce a dynamic maneuvering mechanism for our Multi-Agent System. Just as a rudder provides corrective steering, our system employs a Guard Agent to perform dynamic supervision and intervention. This mechanism steers the Execution Agent’s reasoning process, correcting deviations and ensuring it remains on a robust and accurate path to the final goal. This proactive course-correction is fundamental to achieving the stability and reliability that our architecture provides.
@@
Table 1: In these linearized equations of motion, $u$, $v$, and $r$ represent the velocities in surge (longitudinal), sway (lateral), and yaw, respectively; $\dot{u}$, $\dot{v}$, and $\dot{r}$ are the corresponding linear and angular accelerations. The vessel’s physical properties are given by its mass $m$, its moment of inertia about the vertical axis $I_{Z}$, and the longitudinal position of the center of gravity $x_{G}$. The terms such as $X_{\dot{u}}$, $Y_{v}$, and $N_{r}$ are the hydrodynamic coefficients. Specifically, terms with a dot in the subscript (e.g., $X_{\dot{u}}$, $Y_{\dot{v}}$, $N_{\dot{r}}$) represent added mass and added moment of inertia, while terms without a dot (e.g., $Y_{v}$, $N_{r}$) represent damping coefficients. The constant forward speed of the vessel is $u_{0}$, and $\Delta u$ is the perturbation in surge velocity. Finally, $X_{\text{drift}}$, $Y_{\text{drift}}$, and $N_{\text{drift}}$ represent the external longitudinal force, lateral force, and yaw moment induced by slowly varying wave drift. (Xie et al., 2020)
Figure 3: Zig-Zag tests in regular waves with 90 deg at t = 0 and ratios between wave length and vessel length = 0.75 (Xie et al., 2020)
The core features of the Multi-Agent System (MAS) we developed within the AWorld framework are highlighted in Figure 4. Based on the AWorld framework, we developed a dynamic Multi-Agent System (MAS) that leverages an "agent-as-tool" paradigm and introduces a Guard Agent for rigorous logical verification. Central to our approach is a mechanism of dynamic supervision and maneuvering, which enables adaptive intervention throughout the problem-solving process. Specifically, the Execution Agent initiates the task and dynamically determines when to engage additional agents, guided by system prompts and contextual analysis. Meanwhile, the Guard Agent provides continuous oversight by monitoring, correcting, and issuing reminders about logical reasoning steps, thereby enhancing the Execution Agent’s accuracy and the overall robustness of the solutions. Notably, the Guard Agent is built on the same foundational model as the Execution Agent (e.g., Gemini 2.5 Pro), ensuring greater consistency and improved collaborative capabilities within the system.
Figure 4: Multi-Agent System (MAS) Architecture on AWorld Framework.
@@ 3 Experiments Settings
We compare three distinct methodologies in our experiments. First, the Base approach involves direct question-answering by a single Gemini 2.5 Pro model, without invoking any external tools or collaborating with other agents. Second, the Single Agent System (SAS) pairs a single model (as same as the Base approach: Gemini 2.5 Pro model) with the delicate system prompt and various mcp tools. Here, the model autonomously decides, based on the system prompt, question and context, whether to use external tools or to answer independently. Third, our Multi-Agent System (MAS) extends the SAS setup by introducing the dynamic supervision and maneuvering mechanisms and building a Guard Agent as an additional candidate tool. In this configuration, the Execution Agent can engage the Guard Agent for real-time logical verification during the problem-solving process, fostering more reliable and accurate solutions.
@@ 4 Experimental Results
Gemini 2.5 Pro SAS Gemini 2.5 Pro vs SAS MAS SAS vs MAS
Round 1 Pass@1 32.11% 57.8% 71.56%
Round 2 Pass@1 30.28% 64.22% 65.14%
Round 3 Pass@1 32.11% 65.14% 66.97%
Pass@3 38.53% 81.65% +111.91% 83.49% +2.25%
Pass@1_avg 31.5% 62.39% +98.06% 67.89% +8.82%
Pass@1_std 0.00863 0.03265 +278.33% 0.02701 -17.3%
Table 2: Summary of experimental results across different versions.
Table 2 demonstrates the experimental results across different experimental versions. Introducing the dynamic supervision and maneuvering mechanism and integrating a Guard Agent into the MAS substantially improves problem-solving accuracy. The base Gemini 2.5 Pro model, relying solely on internal knowledge and test-time reasoning, achieves an average pass@1 accuracy of 31.5%. Incorporating tool usage in the SAS nearly doubles this figure, raising the pass@1 score to 62.39% by enabling context expansion and real-world data acquisition. Building on this foundation, the MAS leverages the Guard Agent to calibrate key solution steps, resulting in a further increase to 67.89% pass@1 accuracy—an 8.82% improvement over the SAS, as well as a pass@3 accuracy of 83.49%, a 2.25% gain.
Beyond accuracy, the maneuvering mechanism by integrating a Guard Agent also enhances system stability. While the base model demonstrates a pass@1 standard deviation of 0.0086 at a temperature setting of 0.1, the SAS experiences a nearly fourfold increase in score variance, largely due to uncertainties introduced by external tools. In contrast, the multi-agent setup, thanks to the logical constraints imposed by the Guard Agent, achieves a reduced pass@1 standard deviation of 0.027, a 17.3% reduction compared to the SAS, thereby fostering more consistent performance.
@@ 5 Analysis
A Good Q&A Model Does Not Equal a Good Tool User
The base model (Gemini 2.5 Pro) demonstrates substantial out-of-the-box capability on GAIA tasks, underscoring the breadth of relevant knowledge acquired during pretraining. However, it cannot reliably determine when to rely solely on internal knowledge versus when to invoke external tools for a given problem. Notably, adding tool access does not always preserve internal solution paths; for example, there are tasks that the base model solves in pass@3 that neither the SAS nor the experimental version can solve.
@@ A Good Q&A Model Does Not Equal a Good Tool User
Despite these limitations, experimental results show that tool-integrated agent architectures can dramatically improve accuracy. Such MAS represent promising pathways toward generalized and robust intelligent solutions.
Context Optimization and Logical Convergence
@@
To counteract this instability, we employ a dynamics maneuvering mechanism. Within this architecture, the Execution Agent is enabled to call upon a “Guard Agent” for review whenever it encounters a logical impasse. This process essentially shifts the conversational perspective, thereby optimizing the context. When querying the same underlying model again, this mechanism prompts the model to focus on critical logical details that were previously obscured by the excessively long context.
@@
Experimental data confirms the effectiveness of this method: compared to the SAS, the introduction of the Guard Agent leads to a 17.3% reduction in the pass@1 standard deviation, marking a significant improvement in solution stability and logical consistency.

### v3

# Profile-Aware Maneuvering: A Dynamic Multi-Agent System for Robust GAIA Problem Solving by AWorld

@@ Abstract
The rapid advancement of large language models (LLMs) has empowered intelligent agents to leverage diverse external tools for solving complex real-world problems. However, this reliance introduces new challenges, as extended contexts and noisy tool outputs can undermine system reliability. To address this, we propose a dynamic Multi-Agent System (MAS) in our AWorld framework, where an Execution Agent is supervised by a Guard Agent that provides on-demand "dynamic maneuvering"—verifying and correcting the reasoning process to improve robustness over single-agent systems. To move beyond this generic supervision, we enhance the architecture with a methodology inspired by System Identification from control theory. This method first profiles the Execution Agent offline on a benchmark dataset to create a "performance fingerprint" of its unique weaknesses. The Guard Agent then leverages this fingerprint online to deliver profile-aware supervision, making targeted interventions based on known failure patterns rather than merely reacting to immediate logical flaws. Extensive experiments on the GAIA dataset demonstrate that this profile-aware MAS significantly improves both effectiveness and stability, outperforming not only single-agent systems but also its ’naive’ counterpart. This superior performance led our system to achieve first place among open-source projects on the prestigious GAIA leaderboard. These findings highlight that building truly trustworthy intelligent systems requires not just collaboration, but a deep, empirically-grounded understanding of each agent’s unique capabilities and limitations.
Keywords Dynamic Multi-Agent System  $\cdot$ Runtime  $\cdot$ Stability  $\cdot$ Maneuvering  $\cdot$ System Identification
@@ 1 Introduction
The extraordinary progress of Large Language Models (LLMs) in both capability and scale (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023; The Google DeepMind Team, 2024; Anthropic, 2025) has sparked widespread curiosity about the upper bounds of artificial intelligence. As practitioners push these frontiers, it has become clear that augmenting foundational models with external tools not only expands their problem-solving abilities beyond intrinsic knowledge but also enables the tackling of complex real-world challenges (Kapoor et al., 2024; Huang and Yang, 2025; Krishnan, 2025; Shao et al., 2025). A vivid illustration is the recent IMO competition, where state-of-the-art LLMs struggled in isolation, whereas agent-based systems built upon them solved most tasks (Huang and Yang, 2025). This suggests that the next frontier of AI may lie not only in the power of individual models but also the ingenuity in organizing them for effective collaboration.
@@
This insight has fueled the rapid growth of multi-agent frameworks, yet amid the excitement, a central challenge has emerged: system stability. Empirical results show that agent robustness hinges on the foundational model’s reliability, the nature of integrated tools, and the design of agent orchestration (Coletta et al., 2024; Li et al., 2025; Shojaee et al., 2025). For instance, while the “solver–reviewer” structure is promising, its rigid, dialog-based architecture can introduce extended contexts and practical limitations like fixed turn limits. This underscores a critical need: to advance reliable intelligent systems, research must focus on strategies for building agents that are not just collaborative, but also consistent, resilient, and adaptive.
Drawing inspiration from disciplines like vessel maneuvering—where optimal navigation relies on dynamically adjusting controls rather than using static settings (Xie et al., 2020)—we argue that intelligent agents require a similar principle of dynamic maneuvering. Instead of relying on fixed supervision, agents should adaptively decide when and how to intervene based on the evolving context. To realize this, we constructed a dynamic Multi-Agent System (MAS) within AWorld, our open-source platform. This framework establishes a collaborative dyad: an Execution Agent that performs primary tasks and a Guard Agent that acts as an on-demand supervisor. At critical junctures, the Execution Agent can invoke the Guard Agent to verify reasoning and correct its trajectory, thus implementing our dynamic maneuvering concept.
However, to elevate this system from merely reactive correction to proactive, intelligent guidance, we address its key limitation: a generic Guard Agent, while beneficial, remains blind to its partner’s specific, habitual failure modes and cannot thus provide more targeted supervision and guidance. We overcome this by introducing our core contribution, inspired by System Identification—a mature engineering discipline effectively applied in domains like vessel maneuvering research where understanding dynamic behavior is paramount (Alexandersson et al., 2024; Xue et al., 2021; Xu and Soares, 2013). Our adaptation of this philosophy to model LLM agents bifurcates into two distinct phases. First, in a preparatory offline stage, we systematically benchmark the Execution Agent to generate a detailed "performance fingerprint", a structured profile of its characteristic errors, such as a tendency to hallucinate code. Then, in the crucial online execution stage, the Guard Agent leverages this fingerprint to provide profile-aware maneuvering, actively monitoring for likely failure scenarios and offering targeted, preemptive advice. This transforms the Guard Agent into an expert on its specific partner.
@@
Rigorous testing on the GAIA benchmark (Mialon et al., 2023) demonstrates that our MAS, orchestrated with System Identification, not only surpasses single-agent and naive multi-agent systems but also achieves higher stability as shown in Figure 1. This superior performance led to AWorld achieving first place among open-source projects on the prestigious GAIA test leaderboard, as shown in Figure 2. This research thus highlights the promise of adaptive, profile-aware multi-agent approaches for building powerful and trustworthy intelligent systems equipped to face diverse real-world tasks.
@@ 2 Method
2.1 The Multi-Agent System and Dynamic Maneuvering
Our approach is inspired by well-established principles in control theory, particularly from the field of marine vessel navigation (Xie et al., 2020). The maneuvering of a ship, as modeled by the control equations in Table 1, represents a system subject to complex external forces like wave drift. To ensure the vessel converges to a desired trajectory (e.g., the zig-zag test in Figure 3), continuous, adaptive adjustments are essential. The rudder angle is dynamically controlled not with a fixed setting, but to actively counteract disturbances and guide the ship’s path.
We posit that an agent’s problem-solving process is analogous to this voyage. The agent’s reasoning forms a "trajectory" toward a solution, constantly perturbed by "noise" from long contexts and irrelevant tool outputs—akin to the winds and currents affecting a vessel. This analogy directly informs our foundational system design within the AWorld framework, where we implement a dynamic maneuvering mechanism. As shown in Figure 4, we establish a collaborative dyad:
An Execution Agent, which performs the primary task.
  • An Execution Agent, which performs the primary task.
  • A Guard Agent, which acts as a dynamic supervisor, analogous to a ship’s rudder.
Just as a rudder provides corrective steering, the Execution Agent can invoke the Guard Agent at critical junctures to verify logic and correct deviations, ensuring it remains on a robust path to the final goal.
@@ 2.1 The Multi-Agent System and Dynamic Maneuvering
Table 1: In these linearized equations of motion, terms like $u,v,r$ represent velocities. The key insight for our work lies with the hydrodynamic coefficients (e.g., $X_{\dot{u}}$, $Y_{v}$, $N_{r}$). These coefficients are not universal; they are unique to each ship’s specific hull geometry and mass distribution and must be determined experimentally or numerically. This process of identifying a system’s unique internal parameters from observed behavior is the essence of System Identification. (Alexandersson et al., 2024; Xue et al., 2021; Xu and Soares, 2013)
Figure 3: Phase portrait where the combination of drift angle and yaw rate is shown for zigzag10/10 and zigzag20/20 wPCC model tests. (Alexandersson et al., 2024)
2.2 From Generic Maneuvering to Profile-Aware Guidance
While the foundational MAS architecture improves robustness, its supervision is generic. The Guard Agent reacts to logical errors as they arise but is blind to the Execution Agent’s habitual weaknesses. To achieve a deeper level of intelligence, we draw a more profound inspiration from the practice of System Identification.
2.2.1 Agent Profiling via System Identification
In naval architecture, the equations in Table 1 provide a general model of ship motion. However, to be useful, the specific values of the hydrodynamic coefficients ($X_{\dot{u}},Y_{v}$, etc.) must be determined for each individual vessel. These coefficients, unique to each ship’s hull and design, are identified through a series of standardized experimental tests (like the zig-zag test). This process creates a precise mathematical "fingerprint" of that ship’s unique maneuvering behavior.
We contend that LLM-based agents, like ships, are complex systems with unique, idiosyncratic behaviors. A foundational model provides the general architecture, but how a specific agent instance—shaped by its system prompt and tools—responds to different situations is analogous to a ship’s unknown hydrodynamic coefficients. We therefore adapt the philosophy of System Identification to create an agent profiling methodology as shown in Figure 4. This involves two steps:
  • Systematic Benchmarking: In a preparatory offline stage, we subject the Execution Agent to 62 problems from the GAIA validation dataset (similar distribution but distinct from the test dataset). This is analogous to putting a ship through its sea trials.
  • Fingerprint Generation: We then task a separate, high capacity LLM-based analyzer (Gemini 2.5 Pro in our practice) to study the Execution Agent’s complete input-output logs from the benchmark. The analyzer’s goal is not merely to score performance but to identify and articulate characteristic failure modes. The identification system prompt can be referred in the appendix. It synthesizes these observations into a structured, human readable profile of an agent’s weaknesses—its "performance fingerprint." This fingerprint might detail a tendency to "hallucinate code when a web search fails," a "proclivity for off-by-one errors in array manipulation," or an "over-reliance on the first search result, even if it’s from an unreliable source."
Figure 4: Multi-Agent System built on the Aworld framework: There are three main configurations evaluated in our experiments that can be derived from this figure. Single Agent System: The baseline system, where an Execution Agent uses a set of tools without the Guard Agent. Naive MAS: A Guard Agent is introduced to provide reactive, real time logical verification for the Execution Agent. Profile-Aware MAS: The final system, where the Guard Agent is equipped with a performance fingerprint derived from the System Identification through the offline process, enabling proactive and targeted supervision.
2.2.2 Profile-Aware Supervision in Online Execution
The generated performance fingerprint is the key to transforming the Guard Agent from a naive supervisor into an expert partner. During online task execution, the Guard Agent is "armed" with this fingerprint, typically by injecting it into its system prompt as a core piece of contextual knowledge, as shown in Figure 4.
This fundamentally alters its behavior. Instead of performing generic logical checks, the Guard Agent now provides profile-aware supervision. It actively monitors for scenarios where the Execution Agent is statistically likely to fail according to its profile. For instance, if the fingerprint identifies a weakness in handling multi-step calculations, the Guard Agent will apply heightened scrutiny and may proactively request a step-by-step breakdown whenever such a calculation is initiated. This shift turns its role from reactive correction to targeted, preemptive guidance, directly addressing the agent’s known blind spots and dramatically enhancing the reliability of the entire system.
@@ 3 Experiments Settings
We compare four distinct methodologies in our experiments.
First, the Base approach involves direct question-answering by a single Gemini 2.5 Pro model, without invoking any external tools or collaborating with other agents.
Second, the Single Agent System (SAS) pairs the same foundational model (Gemini 2.5 Pro) with a detailed system prompt and various MCP tools. Here, the model autonomously decides, based on the question and context, whether to use external tools or to answer independently.
Third, our Multi-Agent System (MAS) extends the SAS setup by introducing the dynamic supervision mechanism. This is achieved by building a Guard Agent as an additional candidate tool, which the Execution Agent can engage for real-time logical verification during the problem-solving process. In this configuration, the Guard Agent provides "naive" supervision, as it has no prior knowledge of the Execution Agent’s specific tendencies.
Finally, our Profile-Aware MAS enhances this architecture with our core contribution inspired by System Identification. It builds directly upon the MAS but equips the Guard Agent with a "performance fingerprint" of its partner. This fingerprint is generated in a preparatory offline stage, where the Execution Agent’s behavior is systematically benchmarked on a separate dataset to identify its characteristic failure modes. During the online evaluation, the Guard Agent leverages this fingerprint to provide profile-aware supervision, making targeted interventions based on its partner’s known weaknesses rather than merely reacting to immediate logical inconsistencies.
@@ 4 Experimental Results
LLM SAS SAS vs LLM MAS MAS vs SAS PA-MAS PA-MAS vs MAS
Round 1 Pass@1 32.11% 56.88% 70.64% 72.48%
Round 2 Pass@1 30.28% 63.30% 64.22% 70.64%
Round 3 Pass@1 32.11% 64.22% 66.06% 69.72%
Pass@3 38.53% 80.73% +109.53% 82.57% +2.28% 84.40% +2.22%
Pass@1_avg 31.50% 61.47% +95.14% 66.97% +8.95% 70.95% +5.93%
Pass@1_std 0.00863 0.03265 +279.07% 0.02701 -17.18% 0.01147 -57.41%
Table 2: Summary of experimental results for LLM, SAS, MAS, and Profile-Aware MAS across different rounds.
Table 2 demonstrates the experimental results, showing a progressive improvement across the four methodologies.
The Base model (Gemini 2.5 Pro), which, relying solely on internal knowledge, achieves an average pass@1 accuracy of 31.50%. Introducing tools in the Single Agent System (SAS) nearly doubles this figure, raising the pass@1 to 61.47% by enabling context expansion and real-world data acquisition. Building on this, our Multi-Agent System (MAS) leverages a "naive" Guard Agent for real-time verification, further increasing the pass@1 accuracy to 66.97%, an 8.95% improvement over the SAS. Finally, equipping the Guard Agent with the performance fingerprint in our Profile-Aware MAS pushes these accuracy gains to their peak. This final enhancement further elevates the average pass@1 to 70.95% and the pass@3 to 84.40%. This demonstrates that profile-aware supervision, which proactively targets known failure modes, provides a distinct and measurable advantage over the reactive, generic checks of the naive MAS.
Beyond accuracy, the results reveal the improvement about system stability. While the Base model is highly consistent (standard deviation of 0.00863), the SAS experiences a nearly fourfold increase in score variance, a direct consequence of the uncertainties introduced by external tools. The naive MAS begins to counteract this instability, with the Guard Agent’s logical constraints reducing the standard deviation by 17.18% compared to the SAS. This stabilizing effect becomes dramatically more pronounced with profile-aware maneuvering. The Profile-Aware MAS slashes the standard deviation by an additional 57.41% down to 0.01147, bringing the system’s consistency nearly back to the level of the Base model but at a vastly superior accuracy. This highlights the crucial difference between merely correcting errors after they occur and proactively preventing them based on an empirical understanding of the agent’s behavior.
@@ 5 Analysis
From Recitation to Reasoning: The Mode-Switching Dilemma in LLM Agents
The base model (Gemini 2.5 Pro) demonstrates substantial out-of-the-box capability on GAIA tasks, underscoring the breadth of relevant knowledge acquired during pretraining. However, it cannot reliably determine when to rely solely on internal knowledge versus when to invoke external tools for a given problem. Notably, adding tool access does not always preserve internal solution paths; for example, there are tasks that the base model solves in pass@3 that neither the SAS nor the naive MAS version can solve.
@@
Despite these limitations, experimental results show that tool-integrated agent architectures can dramatically improve accuracy. Such MAS represent promising pathways toward generalized intelligent solutions.
Dynamic Maneuvering: Context Optimization and Logical Convergence
@@
To counteract this instability, we employ a dynamic maneuvering mechanism. Within this architecture, the Execution Agent is enabled to call upon a “Guard Agent” for review whenever it encounters a logical impasse. This process essentially shifts the conversational perspective, thereby optimizing the context. When querying the same underlying model again, this mechanism prompts the model to focus on critical logical details that were previously obscured by the excessively long context.
@@
Experimental data confirms the effectiveness of this method: compared to the SAS, the introduction of the Guard Agent leads to a 17.18% reduction in the pass@1 standard deviation, marking a significant improvement in solution stability and logical consistency.
Beyond Parameter Tuning: System-Level Reinforcement through Explicit Policy
The limitations of reactive supervision highlight a central challenge in AI: how to effectively steer an agent’s behavior toward desired outcomes. The prevailing paradigm for this task is Reinforcement Learning (RL), which fine-tunes a model’s millions of opaque internal parameters based on scalar reward signals. This process implicitly shapes the agent’s decision-making policy, but the policy itself remains a "black box", a complex, distributed state within the model’s weights, making it difficult to interpret, audit, or directly control.
Our work, however, explores a fundamentally different philosophy. Instead of optimizing implicit parameters, our Profile-Aware MAS reinforces the agent’s reasoning process at the system level through an explicit, textual policy. The offline profiling stage acts as a targeted "exploration" phase, but its output is not a simple reward number. Instead, it generates rich, structured feedback that is synthesized into a human-readable "performance fingerprint." During online execution, this fingerprint is injected into the Guard Agent’s prompt, creating what we term Context-Level Reinforcement. The reinforcement signal is not a back-propagated gradient; it is a direct, contextual instruction that guides the reasoning path in real-time.
This distinction is analogous to the difference between subconscious skill acquisition and conscious, expert execution. RL is akin to training an athlete’s muscle memory through repetitive trial and error, where improvement occurs without explicit articulation of the strategy. Our approach, by contrast, is akin to equipping a pilot with a dynamic pre-flight checklist tailored to their known habits and the specific aircraft they are flying. It doesn’t rely on instinct (the internal model) alone, but augments it with conscious, articulated best practices (the explicit fingerprint).
The empirical evidence validates this philosophy. The dramatic results—a further 5.93% gain in pass@1 accuracy and a remarkable 57.41% reduction in standard deviation—demonstrate that this form of explicit, knowledge-based reinforcement is exceptionally effective. It proves that guiding an agent toward more optimal and stable behavior can be achieved more reliably by reinforcing the system with transparent, external knowledge, rather than solely attempting to tune the inscrutable depths of the model itself.
Rethinking Agent Trustworthiness: Beyond Collaboration to Introspection
The distinction between our method and traditional RL extends beyond the mechanism to the core issue of trustworthiness, a cornerstone of AI Safety. Reinforcement Learning, while powerful, often operates as a black-box optimization. The model’s behavior improves, but the "how" and "why" are buried within millions of adjusted numerical weights, rendering the resulting policy opaque and difficult to interpret, audit, or control.
Our framework, by contrast, constitutes a form of engineered introspection. The reinforcement mechanism—the performance fingerprint—is not a set of arcane weights but a human-readable text file. This transparency has profound implications for building dependable systems:
  • We know precisely why the Guard Agent is intervening. Its guidance is based on clear, articulated rules like, "This agent tends to make off-by-one errors; double-check all list indexing."
  • Trustworthiness by Design: Our framework moves towards predictable and auditable AI systems, where reliability is a product of deliberate design, not just empirical performance. Because the corrective logic is transparent, we can build trust not merely because the system works in testing, but because we fundamentally understand how it works. This shifts reliability from an observed property to an engineered characteristic.
  • Controllability & Maintainability: The policy is fully controllable. An engineer can manually review, edit, or augment the fingerprint to refine the system’s behavior without costly retraining. If a new failure mode is discovered in the field, a new rule can be added in minutes, enabling rapid, targeted maintenance.
Ultimately, our findings suggest that the path toward robust AI is not only the implicit tuning of model parameters. By creating methodologies for engineered introspection, we can build systems where reinforcement signals are explicit and the rules of collaboration are themselves a transparent and controllable part of the architecture. This shifts the goal from a quest for better black-box optimizers to the engineering discipline of building truly predictable, reliable, and trustworthy intelligent systems.